OpenStudy (anonymous):

Use the angle sum or difference identity to find the exact value of cos 165 degrees I think I know how to start: cos = 3(45) + 30 Am I on the right track? I'm lost after this part...

OpenStudy (anonymous):

cos=(30+135) use the formula cos(x+y)=cos(x) cos(y)-sin(x)sin(y) Then you would find the individual values of each trig function.

OpenStudy (anonymous):

cos30cos135 - sin30sin135? but why is it minus if it's 30 plus 135?

OpenStudy (anonymous):

Although, I would convert to radians if your problem is in degrees, because I find it is easier to work with radians for questions like these.

OpenStudy (anonymous):

ok I think I worked it out \[(-\sqrt{6}-\sqrt{2})/4\]

OpenStudy (anonymous):

yes? no?

OpenStudy (anonymous):

Sorry, I'm confusing myself. The question asks either the sum or the difference, so you could use either formula, cos (x+y) or cos (x-y). They would probably give the same answer.

OpenStudy (anonymous):

One sec, let me work out the problem and I'll get back to you

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Yes you are correct. Sorry this took so long. I wanted to make sure I was right so I was being very careful and double checked my work.

OpenStudy (anonymous):

no problem. thanks for your help! :)

OpenStudy (anonymous):

np